OpenStudy (anonymous):

Simplify: (x + 5)^2 A. x^2 + 25 B. x^2 + 5x + 25 C. x^2 + 10x + 25 D. x^2 + 10

OpenStudy (anonymous):

(a+b)^2 = a^2 + 2ab + b^2 therefore (x + 5)^2 = x^2+ 10x + 25